OpenStudy (anonymous):

Solve 2x^2 - 20 = 3x.

OpenStudy (anonymous):

THere are many ways my prefer way is to use quadratic formula. Depending on what they want u to use tho they may not like it. BUT 2x2 - 20 = 3x 2x2 - 20 - 3x=0 just subtract 3x from both sides. 2x2 - 3x -20 =0 rearrange now use quadratic equation ax2 + bx +c =0 2x2 - 3x -20

OpenStudy (anonymous):

I mean use quadratic formula

OpenStudy (anonymous):

2 is a -3 is b -20 is c Can you handle that :)

OpenStudy (anonymous):

\[2x^{2}-20=3x\]

OpenStudy (anonymous):

\[2x ^{2}-3x-20=0\]

OpenStudy (anonymous):

\[2x ^{2}-8x+5x-20=0\]

OpenStudy (anonymous):

\[2x(x-4)+5(x-4)=0\] \[\therefore (2x+5)(x-4)=0 \rightarrow x=-2.5 or 4\]